# On the Identification of Temporally Causal Representation with Instantaneous Dependence

### v3

@@ 1 Introduction
Time series analysis (zhang2023self; tang2021probabilistic; li2023difformer; wu2022timesnet; luo2024moderntcn), which has been found widespread applications across diverse fields such as weather (bi2023accurate; wu2023interpretable), finance (tsay2005analysis; huynh2022few), and human activity recognition (yang2015deep; kong2022human), aims to capture the underlying patterns (wang2019time; jin2022multivariate) behind the time series data. To achieve this, one solution is to estimate the latent causal processes (tank2021neural; li2023transferable; zheng2019using). However, without further assumptions (locatello2020sober), it is challenging to identify latent causal processes, i.e., ensure that the estimated latent causal process is correct.
@@
Researchers have exploited Independent Component Analysis (ICA) (hyvarinen2000independent; comon1994independent; hyvarinen2013independent; zhang2007kernel), where observations are generated from latent variables via a linear mixing function, to identify the latent causal process. To deal with nonlinear cases, different types of assumptions including sufficient changes (yao2021learning; khemakhem2020variational; xie2022multi) and structural sparsity (lachapelle2022disentanglement; zheng2022identifiability) are proposed to meet the independent change of latent variables. Specifically, several works leverage auxiliary variables to achieve strong identifiable results of latent variables (halva2020hidden; halva2021disentangling; hyvarinen2016unsupervised; hyvarinen2017nonlinear; khemakhem2020variational; hyvarinen2019nonlinear). To seek identifiability in an unsupervised manner, other researchers (ng2024identifiability; zheng2024generalizing) propose the structural sparsity on the generation process from latent sources to observation. However, these methods usually do not have instantaneous relations, e.g., the latent variables are mutually independent conditional on their time-delayed parents at previous timestamps. Hence, further assumptions are used for identifiability. For example, lippe2023causal demonstrate identifiability by assuming that there exist interventions on latent variables, and morioka2023causal yield identifiability with the assumption of the grouping of observational variables. Recently, zhang2024causal use the sparse structures of latent variables to achieve identifiability of static data under distribution shift. Beyond it, in this work, we demonstrate that temporal identifiability benefits from the sparsity constraint on the causal influences in the latent causal processes by employing the temporal contextual information, i.e., the variables across multiple timestamps. Please refer to Appendix A and D for further discussion of related works, including the identification of latent variables and instantaneous time series.
@@
Figure 1: Three different data generation processes with time-delayed and instantaneous dependencies. (a) iCITRIS (lippe2023causal) requires intervention variables $\textbf{i}_{t}$ for latent variables (the gray nodes with blue rim.). (b) G-CaRL (morioka2023causal) assumes that observed variables can be grouped according to which latent variables they are connected to (the blue dotted lines), (c) IDOL requires a sparse latent causal process (the blue solid lines).
@@
Although these methods (lippe2023causal; morioka2023causal) have demonstrated the identifiability of latent variables with instantaneous dependencies, they often impose strict assumptions on the latent variables and observations. Specifically, as shown in Figure 1 (a), iCITRIS (lippe2023causal) reaches identifiability by assuming that interventions $\bm{I}_{t}$, i.e., the gray nodes with blue rim, on the latent variables, which is expensive and difficult in practice. For instance, considering the joints as the latent variables of a human motion process, it is almost impossible to intervene in the joints of the knee when a human is walking. Moreover, as shown in Figure 1 (b), G-CaRL (morioka2023causal) employs the grouping of observations to identify the latent variables, i.e., the blue dotted lines. However, the grouping of observations is usually expensive, for example, grouping the sensors on a human body from the joints of skeletons requires massive of expert knowledge and labor. Therefore, it is urgent to find a more general approach with an appropriate assumption to identify the latent causal process with instantaneous dependencies.
@@ 2.1 Time Series Generative Model under Instantaneous Latent Causal Process
We begin with the data generation process as shown in Figure 1 (c) where the latent causal process is sparse. To facilitate clarity, we adopt the terminology widely used in ICA literature like observation ${\mathbf{x}}_{t}$, latent variables ${\mathbf{z}}_{t}$ and mixing function $g$ (hyvarinen2016unsupervised). Suppose that we have time series data with discrete timestamps $\textbf{X}=\{{\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t},\cdots,{\mathbf{x}}_{T}\}$, where ${\mathbf{x}}_{t}$ is generated from latent variables ${\mathbf{z}}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{n}$ via an invertible and nonlinear mixing function $\mathbf{g}$}. We have:
@@
Specifically, the $i$-th dimension of the latent variable $z_{t,i}$ is generated via the latent causal process, which is assumed to be related to the time-delayed parent variables $\text{Pa}_{d}(z_{t,i})$ and the instantaneous parents $\text{Pa}_{e}(z_{t,i})$. Formally, it can be written via a structural equation model (neuberg2003causality) as
@@ 2.1.1 Identifiability of Latent Causal Processes
Based on the aforementioned generation process, we further provide the definition of the identifiability of latent causal process with instantaneous dependency in Definition 1. Moreover, if the estimated latent processes can be identified at least up to permutation and component-wise invertible transformation, the latent causal relations are also immediately identifiable up to a Markov equivalence class (spirtes2001causation). We further show how to go beyond the Markov equivalence class and identify the instantaneous causal relations with a mild assumption in Theorem 3.
Definition 1 (Identifiable Latent Causal Process (yao2022temporally; yao2021learning)).
@@ 3 Identification Results for Latent Causal Process
Given the definition of identification of latent causal process, we show how to identify the latent causal process in Figure 1 (c) under a sparse latent process. Please note that our theorem is discussed under Markov Network. For more details, please refer to Appendix B.6. Specifically, we first leverage the connection between conditional independence and cross derivatives (lin1997factorizing) and the sufficient variability of temporal data to discover the relationships between the estimated and the true latent variables, which is shown in Theorem 1. Moreover, we establish the identifiability result of latent variables by enforcing the sparse causal influences, as shown in Theorem 2. Moreover, we also show that the existing identifiability results of the temporally causal representation learning methods are a special case of our IDOL method, which is shown in Corollary 1. Finally, under a mild assumption, we achieve identification of the causal graph underlying the latent causal process, as shown in Theorem 3.
@@ 3.1 Relationships between Ground-truth and Estimated Latent Variables
For a series of observations $\mathbf{x}_{t}\in\mathbb{R}^{n}$ and estimated latent variables $\mathbf{\hat{z}}_{t}\in\mathbb{R}^{n}$ with the corresponding process $\hat{f}_{i},\hat{p}(\epsilon),\hat{g}$, where $\hat{g}$ is invertible, suppose that the process subject to observational equivalence $\mathbf{x}_{t}=\mathbf{\hat{g}}(\mathbf{\hat{z}}_{t})$. Let $\mathbf{c}_{t}\triangleq\{\mathbf{z}_{t-1},\mathbf{z}_{t}\}\in\mathbb{R}^{2n}$ and that $\mathcal{M}_{\mathbf{c}_{t}}$ be the variable set of two consecutive timestamps and the corresponding Markov network respectively. Suppose that the following assumptions hold:
  • A1 (Smooth and Positive Density): The conditional probability function of the latent variables ${\mathbf{c}}_{t}$ is smooth and positive, i.e., $p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})$ is third-order differentiable and $p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})>0$ over $\mathbb{R}^{2n}$,
@@
    $\begin{split}w(m)=&\Big(\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,1}^{2}\partial z_{t-2,m}},\cdots,\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,2n}^{2}\partial z_{t-2,m}}\Big)\oplus\\ &\Big(\frac{\partial^{2}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,1}\partial z_{t-2,m}},\cdots,\frac{\partial^{2}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,2n}\partial z_{t-2,m}}\Big)\oplus\Big(\frac{\partial^{3}\log p({\mathbf{c}}_{t}|{\mathbf{z}}_{t-2})}{\partial c_{t,i}\partial c_{t,j}\partial z_{t-2,m}}\Big)_{(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})},\end{split}$ (5)
@@
    where $\oplus$ denotes concatenation operation and $(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{c}}_{t}})$ denotes all pairwise indice such that $c_{t,i},c_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$. For $m\in[1,\cdots,n]$, there exist $2n+|\mathcal{M}_{{\mathbf{c}}_{t}}|$ different values of ${\mathbf{z}}_{t-2,m}$, such that the $2n+|\mathcal{M}_{{\mathbf{c}}_{t}}|$ values of vector functions $w(m)$ are linearly independent.
Then for any two different entries $\hat{c}_{t,k},\hat{c}_{t,l}$ of $\hat{{\mathbf{c}}}_{t}\in\mathbb{R}^{2n}$ that are not adjacent in the Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ over estimated $\hat{{\mathbf{c}}}_{t}$,
(i) Each ground-truth latent variable $c_{t,i}$ of ${\mathbf{c}}_{t}\in\mathbb{R}^{2n}$ is a function of at most one of $\hat{c}_{k}$ and $\hat{c}_{l}$,
(ii) For each pair of ground-truth latent variables $c_{t,i}$ and $c_{t,j}$ of ${\mathbf{c}}_{t}\in\mathbb{R}^{2n}$ that are adjacent in $\mathcal{M}_{{\mathbf{c}}_{t}}$ over ${\mathbf{c}}_{t}$, they can not be a function of $\hat{c}_{t,k}$ and $\hat{c}_{t,l}$ respectively.
@@
Proof Sketch. The proof can be found in Appendix B.1. First, we establish a bijective transformation between the ground-truth ${\mathbf{z}}_{t}$ and the estimated $\hat{{\mathbf{z}}}_{t}$ to connect them together. Next, we utilize the structural properties of the ground-truth Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$ to constrain the structure of the estimated Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ through the connection between conditional independence and cross derivatives (lin1997factorizing), i.e., $c_{t,i}\perp c_{t,j}|{\mathbf{c}}_{t}\backslash\{c_{t,i},c_{t,j}\}$ implies $\frac{\partial^{2}\log p({\mathbf{c}}_{t})}{\partial c_{t,i}\partial c_{t,j}}=0$. By introducing the sufficient variability assumption, we further construct a linear system with a full rank coefficient matrix to ensure that the only solution holds, i.e.,
$\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,l}}=0,\quad\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,j}}{\partial\hat{c}_{t,l}}=0,\quad\frac{\partial^{2}c_{t,i}}{\partial c_{t,k}\partial c_{t,l}}=0,$ (6)
where $\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,i}}{\partial\hat{c}_{t,l}}=0$ and $\frac{\partial c_{t,i}}{\partial\hat{c}_{t,k}}\cdot\frac{\partial c_{t,j}}{\partial\hat{c}_{t,l}}=0$ correspond to statement (i) statement (ii), respectively.
@@ 3.2 Identification of Latent Variables
In this subsection, we demonstrate that given further sparsity assumptions, the latent Markov network over ${\mathbf{c}}_{t}$ and the latent variables are also identifiable. For a better explanation of the identifiability of latent variables, we first introduce the definition of the Intimate Neighbor Set (zhang2024causal).
Definition 2 (Intimate Neighbor Set (zhang2024causal)).
@@
In other words, $\Psi$ contains all neighbors, iff all neighbors of $c_{t,i}$ are also in the same unique clique as $c_{t,i}$, else $\Psi=\emptyset$. We further provide an example in the Appendix B.8.
Based on the conclusions of Theorem 1, we consider $3$ consecutive timestamps, i.e., ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t},{\mathbf{z}}_{t+1}\}$, where the identifiability of ${\mathbf{z}}_{t}$ can be assured with the help of the contextual information ${\mathbf{z}}_{t-1}$ and ${\mathbf{z}}_{t+1}$.
@@
(Component-wise Identification of Latent Variables with instantaneous dependencies.) Suppose that the observations are generated instantaneous latent process, and $\mathcal{M}_{{\mathbf{c}}_{t}}$ is the Markov network over ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t},{\mathbf{z}}_{t+1}\}\in\mathbb{R}^{3n}$. Except for the smooth, positive density and sufficient variability assumptions, we further make the following assumption:
@@
(i) The estimated Markov network $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$ is isomorphic  [^1] Please refer to the definition of isomorphism of Markov networks in Appendix B.7. to the ground-truth Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$.
@@
Figure 2: Examples of Markov networks with different types of contextual information that satisfy the sparse latent process assumption. Case (i) uses both historical and future information for identifiability. Case (ii) uses only historical information. Case (iii) that does not include instantaneous dependencies can identify latent variables without any contextual information, which degenerates to TDRL (yao2022temporally). Please note that in all cases, we show only the Markov net of ${\mathbf{c}}_{t}$ conditioned on previous timestamps while omitting the conditions for simplicity.
Proof Sketch. Given the fact that there always exists a row permutation for each invertible matrix such that the permuted diagonal entries are non-zero (zheng2024generalizing), we utilize the conclusion of Theorem 1 to demonstrate that any edge present in the true Markov network will necessarily exist in the estimated one. Furthermore, when the number of estimated edges reaches a minimum, the identifiability of Markov network can be achieved. Finally, we once again leverage Theorem 1 to illustrate that the permutation which allows the Markov network to be identified can further lead to a component-wise level identifiability of latent variables. The detailed proof and discussion of assumptions are given in Appendix B.3 and C, respectively.
@@
(General Case for Component-wise Identification.) Suppose that the observations are generated by Equation (1) and (2), and there exists ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-a},\cdots,{\mathbf{z}}_{t},\cdots,{\mathbf{z}}_{t+b}\}\in\mathbb{R}^{(a+b+1)\times n}$ with the corresponding Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$. Suppose assumptions A1 and A2 hold true, and for any $z_{t,i}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{t,i}$ is an empty set. When the observational equivalence is achieved with the minimal number of edges of the estimated Markov network of $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$, there must exist a permutation $\pi$ of the estimated latent variables, such that $z_{t,i}$ and $\hat{z}_{t,\pi(i)}$ is one-to-one corresponding, i.e., $z_{t,i}$ is component-wise identifiable.
Discussion. In this part, we give a further discussion on how contextual information can help to identify current latent variables. Let us take Figure 2 (i) as an example, when we consider the current latent variables $z_{t,1},z_{t,2}$ and $z_{t,3}$ only, it is easy to see that neither $z_{t,1}$ nor $z_{t,3}$ are identifiable since both of their intimate set are $\{z_{t,2}\}$. However, when the historical information ${\mathbf{z}}_{t-1}$ is taken into consideration, where $z_{t-1,1}$ affects only $z_{t,1}$ but not $z_{t,2}$, then $z_{t,1}$ will immediately be identified since the intimate set of $z_{t,1}$ will be empty thanks to the unique influence from $z_{t-1,1}$. Similarly, with a future variable $z_{t+1,3}$ which is influenced by $z_{t,3}$ but not $z_{t,1}$, then $z_{t,3}$ can be identified as well. Besides, under some circumstances, only part of the contextual information is needed. Therefore, the length of the required time window $L$ can be reduced as well, which leads to a more relaxed requirement of assumption A2, according to Theorem 1. As shown in Figure 2 (ii), with only historical information, the identifiability of ${\mathbf{z}}_{t}$ can already be achieved. Furthermore, some existing identifiability results (yao2021learning; yao2022temporally) can be considered as a special case of our theoretical results when the instantaneous effects vanish. For example, Figure 2 (iii) provides the example of TDRL (yao2022temporally), where all intimate sets are empty naturally, and the identifiability is assured.
@@
yao2022temporally $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
morioka2023causal $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
lippe2023causal $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
zhang2024causal $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$

### v4

@@ 1 Introduction
Time series analysis (Zhang et al., 2023; Tang and Matteson, 2021; Li et al., 2023a; Wu et al., 2022; Luo and Wang, 2024), which has been found widespread applications across diverse fields such as weather (Bi et al., 2023; Wu et al., 2023b), finance (Tsay, 2005; Huynh et al., 2022), and human activity recognition (Yang et al., 2015; Kong and Fu, 2022), aims to capture the underlying patterns (Wang et al., 2019; Jin et al., 2022) behind the time series data. To achieve this, one solution is to estimate the latent causal processes (Tank et al., 2021; Li et al., 2023b; Zheng and Kleinberg, 2019). However, without further assumptions (Locatello et al., 2020), it is challenging to identify latent causal processes, i.e., ensure that the estimated latent causal process is correct.
@@
Researchers have exploited Independent Component Analysis (ICA) (Hyvärinen and Oja, 2000; Comon, 1994; Hyvärinen, 2013; Zhang and Chan, 2007), where observations are generated from latent variables via a linear mixing function, to identify the latent causal process. To deal with nonlinear cases, different types of assumptions including sufficient changes (Yao et al., 2021; Khemakhem et al., 2020a; Xie et al., 2022b) and structural sparsity (Lachapelle et al., 2022; Zheng et al., 2022) are proposed to meet the independent change of latent variables. Specifically, several works leverage auxiliary variables to achieve strong identifiable results of latent variables (H"alv"a and Hyvarinen, 2020; Hälvä et al., 2021; Hyvarinen and Morioka, 2016; 2017; Khemakhem et al., 2020a; Hyvarinen et al., 2019). To seek identifiability in an unsupervised manner, other researchers (Ng et al., 2024; Zheng and Zhang, 2024) propose the structural sparsity on the generation process from latent sources to observation. However, these methods usually do not have instantaneous relations, e.g., the latent variables are mutually independent conditional on their time-delayed parents at previous timestamps. Hence, further assumptions are used for identifiability. For example, Lippe et al. (2023) demonstrate identifiability by assuming that there exist interventions on latent variables, and Morioka and Hyvärinen (2023) yield identifiability with the assumption of the grouping of observational variables. Recently, Zhang et al. (2024) use the sparse structures of latent variables to achieve identifiability of static data under distribution shift. Beyond it, in this work, we demonstrate that temporal identifiability benefits from the sparsity constraint on the causal influences in the latent causal processes by employing the temporal contextual information, i.e., the variables across multiple timestamps. Please refer to Appendix A and D for further discussion of related works, including the identification of latent variables and instantaneous time series.
@@
Figure 1: Three different data generation processes with time-delayed and instantaneous dependencies. (a) iCITRIS (Lippe et al., 2023) requires intervention variables $\textbf{i}_{t}$ for latent variables (the gray nodes with blue rim.). (b) G-CaRL (Morioka and Hyvärinen, 2023) assumes that observed variables can be grouped according to which latent variables they are connected to (the blue dotted lines), (c) IDOL requires a sparse latent causal process (the blue solid lines).
@@
Although these methods (Lippe et al., 2023; Morioka and Hyvärinen, 2023) have demonstrated the identifiability of latent variables with instantaneous dependencies, they often impose strict assumptions on the latent variables and observations. Specifically, as shown in Figure 1 (a), iCITRIS (Lippe et al., 2023) reaches identifiability by assuming that interventions $\bm{I}_{t}$, i.e., the gray nodes with blue rim, on the latent variables, which is expensive and difficult in practice. For instance, considering the joints as the latent variables of a human motion process, it is almost impossible to intervene in the joints of the knee when a human is walking. Moreover, as shown in Figure 1 (b), G-CaRL (Morioka and Hyvärinen, 2023) employs the grouping of observations to identify the latent variables, i.e., the blue dotted lines. However, the grouping of observations is usually expensive, for example, grouping the sensors on a human body from the joints of skeletons requires massive of expert knowledge and labor. Therefore, it is urgent to find a more general approach with an appropriate assumption to identify the latent causal process with instantaneous dependencies.
@@ 2.1 Time Series Generative Model under Instantaneous Latent Causal Process
We begin with the data generation process as shown in Figure 1 (c) where the latent causal process is sparse. To facilitate clarity, we adopt the terminology widely used in ICA literature like observation ${\mathbf{x}}_{t}$, latent variables ${\mathbf{z}}_{t}$ and mixing function $g$ (Hyvarinen and Morioka, 2016). Suppose that we have time series data with discrete timestamps $\textbf{X}=\{{\mathbf{x}}_{1},\cdots,{\mathbf{x}}_{t},\cdots,{\mathbf{x}}_{T}\}$, where ${\mathbf{x}}_{t}$ is generated from latent variables ${\mathbf{z}}_{t}\in\mathcal{Z}\subseteq\mathbb{R}^{n}$ via an invertible and nonlinear mixing function $\mathbf{g}$}. We have:
@@
Specifically, the $i$-th dimension of the latent variable $z_{t,i}$ is generated via the latent causal process, which is assumed to be related to the time-delayed parent variables $\text{Pa}_{d}(z_{t,i})$ and the instantaneous parents $\text{Pa}_{e}(z_{t,i})$. Formally, it can be written via a structural equation model (Neuberg, 2003) as
@@ 2.1.1 Identifiability of Latent Causal Processes
Based on the aforementioned generation process, we further provide the definition of the identifiability of latent causal process with instantaneous dependency in Definition 1. Moreover, if the estimated latent processes can be identified at least up to permutation and component-wise invertible transformation, the latent causal relations are also immediately identifiable up to a Markov equivalence class (Spirtes et al., 2001). We further show how to go beyond the Markov equivalence class and identify the instantaneous causal relations with a mild assumption in Theorem 3.
Definition 1 (Identifiable Latent Causal Process (Yao et al., 2022; 2021)).
@@ 3 Identification Results for Latent Causal Process
Given the definition of identification of latent causal process, we show how to identify the latent causal process in Figure 1 (c) under a sparse latent process. Please note that our theorem is discussed under Markov Network. For more details, please refer to Appendix B.6. Specifically, we first leverage the connection between conditional independence and cross derivatives (Lin, 1997) and the sufficient variability of temporal data to discover the relationships between the estimated and the true latent variables, which is shown in Theorem 1. Moreover, we establish the identifiability result of latent variables by enforcing the sparse causal influences, as shown in Theorem 2. Moreover, we also show that the existing identifiability results of the temporally causal representation learning methods are a special case of our IDOL method, which is shown in Corollary 1. Finally, under a mild assumption, we achieve identification of the causal graph underlying the latent causal process, as shown in Theorem 3.
@@ 3.1 Relationships between Ground-truth and Estimated Latent Variables
  • A2 (Sufficient Variability): Denote $|\mathcal{M}_{{\mathbf{z}}_{t}}|$ as the number of edges in Markov network $\mathcal{M}_{{\mathbf{z}}_{t}}$. Let
@@
    $\begin{split}w(m)=&\Big(\frac{\partial^{3}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,1}^{2}\partial z_{t-1,m}},\cdots,\frac{\partial^{3}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,n}^{2}\partial z_{t-1,m}}\Big)\oplus\\ &\Big(\frac{\partial^{2}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,1}\partial z_{t-1,m}},\cdots,\frac{\partial^{2}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,n}\partial z_{t-1,m}}\Big)\oplus\Big(\frac{\partial^{3}\log p({\mathbf{z}}_{t}|{\mathbf{z}}_{t-1})}{\partial z_{t,i}\partial z_{t,j}\partial z_{t-1,m}}\Big)_{(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}_{t}})},\end{split}$ (5)
@@
    where $\oplus$ denotes concatenation operation and $(i,j)\in\mathcal{E}(\mathcal{M}_{{\mathbf{z}}_{t}})$ denotes all pairwise indice such that $z_{t,i},z_{t,j}$ are adjacent in $\mathcal{M}_{{\mathbf{z}}_{t}}$. For $m\in[1,\cdots,n]$, there exist $2n+|\mathcal{M}_{{\mathbf{z}}_{t}}|$ different values of ${\mathbf{z}}_{t-1,m}$, such that the $2n+|\mathcal{M}_{{\mathbf{z}}_{t}}|$ values of vector functions $w(m)$ are linearly independent.
@@
Proof Sketch. The proof can be found in Appendix B.1. First, we establish a bijective transformation between the ground-truth ${\mathbf{z}}_{t}$ and the estimated $\hat{{\mathbf{z}}}_{t}$ to connect them together. Next, we utilize the structural properties of the ground-truth Markov network $\mathcal{M}_{{\mathbf{z}}_{t}}$ to constrain the structure of the estimated Markov network $\mathcal{M}_{\hat{{\mathbf{z}}}_{t}}$ through the connection between conditional independence and cross derivatives (Lin, 1997), i.e., $z_{t,i}\perp z_{t,j}|{\mathbf{z}}_{t}\backslash\{z_{t,i},z_{t,j}\}$ implies $\frac{\partial^{2}\log p({\mathbf{z}}_{t})}{\partial z_{t,i}\partial z_{t,j}}=0$. By introducing the sufficient variability assumption, we further construct a linear system with a full rank coefficient matrix to ensure that the only solution holds, i.e.,
$\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,l}}=0,\quad\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,l}}=0,\quad\frac{\partial^{2}z_{t,i}}{\partial z_{t,k}\partial z_{t,l}}=0,$ (6)
where $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,i}}{\partial\hat{z}_{t,l}}=0$ and $\frac{\partial z_{t,i}}{\partial\hat{z}_{t,k}}\cdot\frac{\partial z_{t,j}}{\partial\hat{z}_{t,l}}=0$ correspond to statement (i) statement (ii), respectively.
@@ 3.2 Identification of Latent Variables
In this subsection, we demonstrate that given further sparsity assumptions, the latent Markov network over ${\mathbf{z}}_{t}$ and the latent variables are also identifiable. For a better explanation of the identifiability of latent variables, we first introduce the definition of the Intimate Neighbor Set (Zhang et al., 2024).
Definition 2 (Intimate Neighbor Set (Zhang et al., 2024)).
@@
In other words, $\Psi$ contains all neighbors, iff all neighbors of $z_{t,i}$ are also in the same unique clique as $z_{t,i}$, else $\Psi=\emptyset$. We further provide an example in the Appendix B.8.
Based on the conclusions of Theorem 1, we consider $3$ consecutive timestamps, i.e., ${\mathbf{z}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t}\}$, where the identifiability of ${\mathbf{z}}_{t}$ can be assured with the help of the contextual information ${\mathbf{z}}_{t-1}$ and ${\mathbf{z}}_{t+1}$.
@@
(Component-wise Identification of Latent Variables with instantaneous dependencies.) Suppose that the observations are generated instantaneous latent process, and $\mathcal{M}_{{\mathbf{z}}_{t}}$ is the Markov network over ${\mathbf{z}}_{t}=\{{\mathbf{z}}_{t-1},{\mathbf{z}}_{t},{\mathbf{z}}_{t+1}\}\in\mathbb{R}^{2n}$. Except for the smooth, positive density and sufficient variability assumptions, we further make the following assumption:
@@
When the observational equivalence is achieved with the minimal number of edges of the estimated Markov network of $\mathcal{M}_{\hat{{\mathbf{z}}}_{t}}$, then we have the following two statements:
(i) The estimated Markov network $\mathcal{M}_{\hat{{\mathbf{z}}}_{t}}$ is isomorphic  [^1] Please refer to the definition of isomorphism of Markov networks in Appendix B.7. to the ground-truth Markov network $\mathcal{M}_{{\mathbf{z}}_{t}}$.
@@
Figure 2: Examples of Markov networks with different types of contextual information that satisfy the sparse latent process assumption. Case (i) uses both historical and future information for identifiability. Case (ii) uses only historical information. Case (iii) that does not include instantaneous dependencies can identify latent variables without any contextual information, which degenerates to TDRL (Yao et al., 2022). Please note that in all cases, we show only the Markov net of ${\mathbf{z}}_{t}$ conditioned on previous timestamps while omitting the conditions for simplicity.
Proof Sketch. Given the fact that there always exists a row permutation for each invertible matrix such that the permuted diagonal entries are non-zero (Zheng and Zhang, 2024), we utilize the conclusion of Theorem 1 to demonstrate that any edge present in the true Markov network will necessarily exist in the estimated one. Furthermore, when the number of estimated edges reaches a minimum, the identifiability of Markov network can be achieved. Finally, we once again leverage Theorem 1 to illustrate that the permutation which allows the Markov network to be identified can further lead to a component-wise level identifiability of latent variables. The detailed proof and discussion of assumptions are given in Appendix B.3 and C, respectively.
@@
(General Case for Component-wise Identification.) Suppose that the observations are generated by Equation (1) and (2), and there exists ${\mathbf{c}}_{t}=\{{\mathbf{z}}_{t-a},\cdots,{\mathbf{z}}_{t}\}$ with the corresponding Markov network $\mathcal{M}_{{\mathbf{c}}_{t}}$. Suppose assumptions A1 and A2 hold true, and for any $z_{t,i}\in{\mathbf{z}}_{t}$, the intimate neighbor set of $z_{t,i}$ is an empty set. When the observational equivalence is achieved with the minimal number of edges of the estimated Markov network of $\mathcal{M}_{\hat{{\mathbf{c}}}_{t}}$, there must exist a permutation $\pi$ of the estimated latent variables, such that $z_{t,i}$ and $\hat{z}_{t,\pi(i)}$ is one-to-one corresponding, i.e., $z_{t,i}$ is component-wise identifiable.
Discussion. In this part, we give a further discussion on how contextual information can help to identify current latent variables. Let us take Figure 2 (i) as an example, when we consider the current latent variables $z_{t,1},z_{t,2}$ and $z_{t,3}$ only, it is easy to see that neither $z_{t,1}$ nor $z_{t,3}$ are identifiable since both of their intimate set are $\{z_{t,2}\}$. However, when the historical information ${\mathbf{z}}_{t-1}$ is taken into consideration, where $z_{t-1,1}$ affects only $z_{t,1}$ but not $z_{t,2}$, then $z_{t,1}$ will immediately be identified since the intimate set of $z_{t,1}$ will be empty thanks to the unique influence from $z_{t-1,1}$. Similarly, with a future variable $z_{t+1,3}$ which is influenced by $z_{t,3}$ but not $z_{t,1}$, then $z_{t,3}$ can be identified as well. Besides, under some circumstances, only part of the contextual information is needed. Therefore, the length of the required time window $L$ can be reduced as well, which leads to a more relaxed requirement of assumption A2, according to Theorem 1. As shown in Figure 2 (ii), with only historical information, the identifiability of ${\mathbf{z}}_{t}$ can already be achieved. Furthermore, some existing identifiability results (Yao et al., 2021; 2022) can be considered as a special case of our theoretical results when the instantaneous effects vanish. For example, Figure 2 (iii) provides the example of TDRL (Yao et al., 2022), where all intimate sets are empty naturally, and the identifiability is assured.
@@
Yao et al. (2022) $\checkmark$ $\checkmark$ $\checkmark$ $\times$ $\checkmark$
Morioka and Hyvärinen (2023) $\checkmark$ $\times$ $\checkmark$ $\checkmark$ $\checkmark$
Lippe et al. (2023) $\times$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$
Zhang et al. (2024) $\checkmark$ $\checkmark$ $\times$ $\checkmark$ $\times$